The sum of final and initial velocity is 17. Acceleration is 10m/s^2 and displacement is .85 m. Find initial and final velocity.

Grade:10

2 Answers

Lakshmi Warrier
6 years ago
From my observation, i think time is also neede to answer this qn...
By the way, is this an mcq?
Given,
v+u=17 -----(1)
(v-u)/t=10ms-2
v-u=10ms-2 * t seconds
     = 10t m/s  ----- (2)
 Solve 1 and 2
v=(17+10t)/2
Substitute in (1)
(17+10t)/2 +u = 17
17+10t+2u= 34
10t + 2u=17
u = (17- 10t)/2
 Substitue in (1)
 v+17+10t =17
v+10t=0
v= -10t
this is why i said if t was there, answer would be accurate
 
 
and if anyone knows it more perfect, pls write that too
 
anjace
6 years ago
we know that v+u=17;
u=17-v
using 3 eq of motion
v^2-u^2=2as
by putting value of v in this equation we get v
therefore, v=9and u=8; thnx if it is right tell me